# DeepProofLog: Efficient Proving in Deep Stochastic Logic Programs

### v1

@@ Abstract
Neurosymbolic (NeSy) AI aims to combine the strengths of neural architectures and symbolic reasoning to improve the accuracy, interpretability, and generalization capability of AI models. While logic inference on top of subsymbolic modules has been shown to effectively guarantee these properties, this often comes at the cost of reduced scalability, which can severely limit the usability of NeSy models. This paper introduces DeepProofLog (DPrL), a novel NeSy system based on stochastic logic programs, which addresses the scalability limitations of previous methods. DPrL parameterizes all derivation steps with neural networks, allowing efficient neural guidance over the proving system. Additionally, we establish a formal mapping between the resolution process of our deep stochastic logic programs and Markov Decision Processes, enabling the application of dynamic programming and reinforcement learning techniques for efficient inference and learning. This theoretical connection improves scalability for complex proof spaces and large knowledge bases. Our experiments on standard NeSy benchmarks and knowledge graph reasoning tasks demonstrate that DPrL outperforms existing state-of-the-art NeSy systems, advancing scalability to larger and more complex settings than previously possible.
@@ 1 Introduction
Neurosymbolic (NeSy) AI represents a promising paradigm that seeks to bridge the gap between data-driven neural architectures and structured reasoning capabilities of symbolic methods to obtain improved accuracy and enhanced interpretability. A key research direction within NeSy focuses on integrating probability with logic programming, enabling principled learning and inference under uncertainty (manhaeve2018deepproblog; dai2019bridging; marra2021neural; pryor2022neupsl; maene2023soft; yang2023neurasp; de2023neural). These systems typically perform inference in two stages: (1) apply symbolic proving to find proofs for a given query, (2) rely on probabilistic inference over these proofs to compute the probability that the query holds. This process employs the possible world semantics (de2007problog; manhaeve2018deepproblog), which computes the probability of a query as the sum of the probabilities of those possible worlds that have at least a proof for the query. Regrettably, this probabilistic inference task corresponds to the weighted model counting problem, which is #P-hard (roth1996hardness).
To alleviate the scalability challenges in probabilistic-logic NeSy systems (wang2013programming; maene2024hardness; marra2024statistical; jiao2024valid), most existing studies focus on scaling the probabilistic inference phase, assuming that derivations of queries can be computed tractably, including sampling approaches (verreet2024explain) and variational approaches (abboud2020learning; dos2021neural; van2023nesi). Although these methods represent progress in the scalability of NeSy, the applicability to large-scale relational settings remains out of their reach. This limitation arises from two primary factors. First, many solutions are still ad hoc, tailored to specific tasks, and require substantial manual effort to adapt to new domains. Second, and more fundamentally, these methods do not address the computational cost of deriving the proofs.
@@
This paper introduces DeepProofLog (DPrL), a novel system taking a different perspective to significantly advance NeSy scalability (see Figure 1). Instead of relying on possible world semantics, our work focuses on the semantics of Stochastic Logic Programs (SLPs) (cussens2001parameter), which define a probability distribution over derivations rather than over possible worlds. The probability associated with a clause in SLPs reflects its likelihood of being used in a successful query derivation, rather than its truth in a possible world. In this perspective, SLPs do not directly model uncertainty about the world, but rather the uncertainty in the behavior of a prover operating within that world. This view offers an agent-based interpretation of inference and learning in NeSy, which, to the best of our knowledge, has not been previously explored. Furthermore, we define a formal correspondence between learning and inference in (deep) SLPs and the problem of learning optimal policies in Markov Decision Processes (MDPs) (feinberg2012handbook). In our formulation, each derivation corresponds to a trajectory in the MDP, and each proving step is parameterized by a neural network that works as a policy by providing continuous and goal-directed guidance for selecting the next action (i.e. a rule application or variable substitution). This connection allows us to draw upon the rich body of established MDP solution methods, which have proven effective in a variety of complex domains, such as game playing and robotics, to tackle scenarios that have traditionally been intractable for NeSy systems. In particular, when the symbolic goal space is computationally manageable, we show that dynamic programming techniques can be used for exact inference with enhanced efficiency. Conversely, when the complete goal space becomes intractable, we apply deep reinforcement learning to approximate the optimal policy through learning from sampled derivations.
@@
To our knowledge, the only other NeSy system designed as a neural extension of SLPs is DeepStochLog (winters2022deepstochlog). It is based on stochastic definite clause grammars, a type of SLPs, and introduces neural grammar rules. While DeepStochLog represents a step forward, it still faces challenges with scalability and expressiveness compared to DPrL. A more technical comparison between our method and DeepStochLog is reported in Section 7.
@@
Contributions. This paper makes the following contributions. (i) We introduce DeepProofLog (DPrL), a novel deep SLP model that leverages neural parameterization at each proof step to provide continuous, goal-directed guidance during reasoning. (ii) We propose a theoretical mapping between derivation-based probabilistic logic reasoning and MDPs, enabling the use of powerful MDP solution techniques for efficient learning and inference in DPrL. (iii) The experiments on classic NeSy benchmarks and knowledge graph completion tasks empirically validate the effectiveness of our approach, achieving improved task performance, efficiency, and scalability with respect to other NeSy systems.
@@ 2.1 Logic Programming
Following (flach1994simply; lloyd2012foundations), we summarize the basic concepts of logic programming. A term is either a constant, a variable, or a compound term of the form $f(t_{1},...,t_{n})$, where $f$ is a functor of arity $n$ and each $t_{i}$ is a term. An expression $r(t_{1},...,t_{n})$, where $r$ is a predicate and each $t_{i}$ is a term, is called an atom. For example, $\textit{locatedIn}(X,\textit{europe})$ is an atom where europe denotes a constant and locatedIn a binary predicate. A definite clause [^2] We will omit “definite” as all clauses in this paper are definite. is of the form $H\leftarrow B_{1},...,B_{m}$, where $H$ is the head atom and $B_{1},\dots,B_{m}$ form the body atoms. A set of definite clauses $\mathcal{P}$ is called a definite logic program. A clause of the form $\leftarrow A_{1},\dots,A_{n}$ is called goal and represents a query to the program. A substitution $\theta$ such as $\{X/italy\}$ is a mapping from variables to terms. A substitution $\theta$ is a unifier of two atoms (same for terms) $A_{1}$ and $A_{2}$ if $A_{1}\theta$ (application of $\theta$ to variables of $A_{1}$) = $A_{2}\theta$. For instance, $\theta=\{X/\textit{italy},Y/\textit{europe}\}$ unifies $A_{1}=\textit{locatedIn}(X,\textit{europe})$ and $A_{2}=\textit{locatedIn}(\textit{italy},Y)$, as $A_{1}\theta=A_{2}\theta=\textit{locatedIn}(\textit{italy},\textit{europe})$. A substitution $\theta$ is the Most General Unifier (MGU) if any other unifier can be obtained by further instantiating $\theta$—i.e., it makes the minimal necessary variable substitutions.
@@ 2.2 Stochastic SLD Resolution
Definition 1 (Probability Distribution over SLD Derivations).
Let $d=(G_{0},G_{1},\dots,G_{n})$ be a finite SLD derivation. The probability of $d$ is defined as:
@@
Let $\mathcal{R}(q)$ be the set of all successful derivations for a query $q$, i.e. derivations $d=(G_{0},G_{1},...,G_{n})$ with $G_{0}=q$ and $G_{n}=\textit{True}$, for some $n>0$. The success probability of $q$ is:
@@
DeepStochLog. DeepStochLog replaces fixed clause probabilities of SLPs with a neural network $\rho_{\lambda}$, which produces a probability distribution over possible variable substitutions. This induces a term-level neural parameterization of the transition model $p(G_{i+1}\!\mid\!G_{i})\!=\!\rho_{\lambda}(X\theta_{i},Y\theta_{i})$, where $\theta_{i}$ is the MGU used in the resolution step from $G_{i}$ to $G_{i+1}$, i.e. $\rho_{\lambda}$ is conditioned on the instantiated input and output terms $X\theta_{i},Y\theta_{i}$ realizing the unification step. An example is shown in Figure 1.
@@ 2.3 Markov Decision Processes
An MDP is a formal model of decision-making defined by the tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is a set of states, $\mathcal{A}(s)$ is a set of actions available in state $s$, $P(s,a,s^{\prime})$ is the transition function giving the probability of transitioning from state $s$ to $s^{\prime}$ with action $a$, $R(s,a,s^{\prime})$ is the reward function giving the expected reward for the transition and $\gamma\in[0,1]$ is the discount factor for future rewards. A policy $\pi(a|s)$ defines the probability of selecting action $a$ in state $s$. Given a starting state $s_{0}$, solving an MDP requires finding the optimal policy $\pi^{*}$ that maximizes the future cumulative reward, defined as the value function:
@@ 3 DeepProofLog
To address the limitations of previous approaches, we propose a shift in perspective: instead of parameterizing the logic program itself (e.g., via clause weights), we parameterize the resolution process. To this aim we introduce DeepProofLog, which defines clause selection probabilities conditioned on the entire current goal and enables dynamic adaptation to vary substitution spaces. DeepProofLog is a framework for differentiable, goal-conditioned SLD resolution that supports flexible and context-aware reasoning.
@@
where $\mathcal{N}(G)$ denotes the set of next goals from resolving the leftmost atom of $G$; $f_{\lambda}(G,G^{\prime})$ scores the compatibility between the learned embeddings of the current goal $G$ and candidate next goal $G^{\prime}$ (as defined in the next paragraph); and $\lambda$ comprises all model parameters.
@@
(Sub)symbolic Embeddings. Let $\mathcal{V}$ denote the set of all entities involved in reasoning, including symbolic tokens (e.g., predicates, constants, variables) and subsymbolic inputs (e.g., images). Each element $v\in\mathcal{V}$ is associated with a learnable dense embedding $\mathbf{e}_{v}\in\mathbb{R}^{d}$. The embeddings $e_{v}$ for all $v\in\mathcal{V}$ are part of $\lambda$ (cf. Equation 1) and trained jointly with the reasoning model. This design enables a unified embedding space in which both symbolic and subsymbolic information can interact seamlessly during inference and learning.
@@
Compatibility Score. The compatibility score between goals $G$ and $G^{\prime}$ is computed as the scalar product between their embeddings: $f_{\lambda}(G,G^{\prime})=\mathbf{e}_{G}\cdot\mathbf{e}_{G^{\prime}}$.
@@
In line with similar approaches (cf. Section 2), DeepProofLog defines the query success probability like in Stochastic SLD resolution (cf. Definition 2). Moreover, our learning objective aligns with DeepStochLog,
@@ 4 Markov Decision Processes for DeepProofLog
DeepProofLog’s novel perspective of parameterizing stochastic SLD resolution opens the door to interpreting learning and reasoning in any (deep) SLP as learning an optimal policy within a logic program-based environment of an MDP. In the following we show how DeepProofLog instantiates an MDP by defining the components $(\mathcal{S},\mathcal{A},P,R,\gamma)$.
@@
Reward. To define the reward, we introduce an auxiliary set of labeled states, where each state is paired with the label of the original query it stems from, i.e., $\tilde{\mathcal{S}}=\{(G,y^{(i)})\mid G\in\mathcal{S},y^{(i)}\in\{0,1\}\text{ is the label of }q^{(i)}\in\mathcal{Q}\}$. Note that the labeled states $\tilde{\mathcal{S}}$ are only used for reward computation, and they do not affect the state space $\mathcal{S}$, and are not accessed by the policy. The reward function $R:\tilde{\mathcal{S}}\to\mathbb{R}$ is defined as
@@
As in standard MDPs, the value function of a policy is defined as its expected cumulative reward. While value functions generally follow the Bellman recursion, our SLD resolution MDP allows a simplified formulation for the value of a non-terminal goal $G$ with label $y$ (see Appendix A.2 for derivation):
@@
We illustrate the introduced terms with the following example. Consider a logic program $\mathcal{P}=\{C_{1},C_{2},C_{3},C_{4},C_{5}\}$, where the clauses are defined as: $C_{1}=\textit{locIn}(X,Y)\leftarrow\textit{neighOf}(X,Z),\textit{locIn}(Z,Y)$, $C_{2}={}\leftarrow\textit{neighOf}(\textit{it},\textit{fr})$, $C_{3}={}\leftarrow\textit{locIn}(\textit{fr},\textit{eu})$, $C_{4}={}\leftarrow\textit{locIn}(\textit{tr},\textit{gr})$, and $C_{5}={}\leftarrow\textit{locIn}(\textit{gr},\textit{eu})$. Consider a positive query $q^{(0)}=G_{0}^{(0)}={}\leftarrow\textit{locIn}(\textit{it},\textit{eu})$ with label $y^{(0)}=1$. The corresponding action space is $\mathcal{A}(G_{0}^{(0)})=\{(C_{1},\{X/\textit{it},Z/\textit{eu}\})\}$. Applying the action gives the next goal $G_{1}^{(0)}={}\leftarrow\textit{neighOf}(\textit{it},Y),\textit{locIn}(Y,\textit{eu})$, with intermediate reward $R(G_{1}^{(0)})=0$. Assuming the episode ends after $n$ resolution steps with $G_{n}^{(0)}=\textit{True}$, the final reward is $R(G_{n}^{(0)},y^{(0)})=2\cdot 1-1=1$. Now consider a negative query $q^{(1)}=G_{0}^{(1)}={}\leftarrow\textit{locIn}(\textit{tr},\textit{eu})$ with label $y^{(1)}=0$. If the episode ends in $G_{m}^{(1)}=\textit{True}$ after $m$ resolution steps, the final reward is $R(G_{m}^{(1)},y^{(1)})=2\cdot 0-1=-1$.

### v2

@@ Abstract
Neurosymbolic (NeSy) AI combines neural architectures and symbolic reasoning to improve accuracy, interpretability, and generalization. While logic inference on top of subsymbolic modules has been shown to effectively guarantee these properties, this often comes at the cost of reduced scalability, which can severely limit the usability of NeSy models. This paper introduces DeepProofLog (DPrL), a novel NeSy system based on stochastic logic programs, which addresses the scalability limitations of previous methods. DPrL parameterizes all derivation steps with neural networks, allowing efficient neural guidance over the proving system. Additionally, we establish a formal mapping between the resolution process of our deep stochastic logic programs and Markov Decision Processes, enabling the application of dynamic programming and reinforcement learning techniques for efficient inference and learning. This theoretical connection improves scalability for complex proof spaces and large knowledge bases. Our experiments on standard NeSy benchmarks and knowledge graph reasoning tasks demonstrate that DPrL outperforms existing state-of-the-art NeSy systems, advancing scalability to larger and more complex settings than previously possible.
Codehttps://github.com/DeepProofLog/DPrL-AAAI
@@ 1 Introduction
Neurosymbolic (NeSy) AI represents a promising paradigm that seeks to bridge the gap between data-driven neural architectures and the structured reasoning capabilities of symbolic methods to obtain improved accuracy and enhanced interpretability. A key research direction within NeSy focuses on integrating probability with logic programming, enabling principled learning and inference under uncertainty (Manhaeve et al. 2018; Dai et al. 2019; Marra and Kuželka 2021; Pryor et al. 2022; Maene and De Raedt 2023; Yang et al. 2023; De Smet et al. 2023). These systems typically perform inference in two stages: (1) apply symbolic proving to find proofs for a given query, (2) rely on probabilistic inference over these proofs to compute the probability that the query holds. This process employs the possible world semantics (De Raedt et al. 2007; Manhaeve et al. 2018), which computes the probability of a query as the sum of the probabilities of those possible worlds that have at least one proof for the query. Regrettably, this probabilistic inference task corresponds to the weighted model counting problem, which is #P-hard (Roth 1996).
To alleviate the scalability challenges in probabilistic-logic NeSy systems (Wang et al. 2013; Maene et al. 2024; Marra et al. 2024; Jiao et al. 2024), most existing studies focus on scaling the probabilistic inference phase, assuming that derivations of queries can be computed tractably, including sampling approaches (Verreet et al. 2024) and variational approaches (Abboud et al. 2020; Dos Martires 2021; van Krieken et al. 2023). Although these methods represent progress in the scalability of NeSy, their applicability to large-scale relational settings remains out of reach. This limitation arises from two primary factors. First, many solutions are still ad hoc, tailored to specific tasks, and require substantial manual effort to adapt to new domains. Second, and more fundamentally, these methods do not address the computational cost of deriving the proofs.
@@
This paper introduces DeepProofLog (DPrL), a novel system that takes a different perspective to significantly advance NeSy scalability (see Figure 1). Instead of relying on possible world semantics, our work focuses on the semantics of Stochastic Logic Programs (SLPs) (Cussens 2001), which define a probability distribution over derivations rather than over possible worlds. The probability associated with a clause in SLPs reflects its likelihood of being used in a successful query derivation rather than its truth in a possible world. From this perspective, SLPs do not directly model uncertainty about the world, but rather the uncertainty in the behavior of a prover operating within that world. This view offers an agent-based interpretation of inference and learning in NeSy, which, to the best of our knowledge, has not been previously explored. Furthermore, we define a formal correspondence between learning and inference in (deep) SLPs and the problem of learning optimal policies in Markov Decision Processes (MDPs) (Feinberg and Shwartz 2012). In our formulation, each derivation corresponds to a trajectory in the MDP, and each proving step is parameterized by a neural network that works as a policy by providing continuous and goal-directed guidance for selecting the next action (i.e. a rule application or variable substitution). This connection allows us to draw upon the rich body of established MDP solution methods, which have proven effective in a variety of complex domains, such as game playing and robotics, to tackle scenarios that have traditionally been intractable for NeSy systems. In particular, when the symbolic goal space is computationally manageable, we show that dynamic programming techniques can be used for exact inference with enhanced efficiency. Conversely, when the complete goal space becomes intractable, we apply deep reinforcement learning (RL) to approximate the optimal policy through learning from sampled derivations.
While automated theorem proving (ATP) can also be framed as an MDP and solved with RL, NeSy reasoning differs conceptually. ATP aims to find any valid proof, whereas NeSy focuses on probabilistic logics, weighing multiple proofs and assigning low probabilities to those inconsistent with subsymbolic signals such as images or embeddings. Consequently, DPrL must both identify plausible proofs and ensure their probabilities align with these subsymbolic signals—this alignment is the central challenge we address to improve efficiency in NeSy learning.
@@
To our knowledge, the only other NeSy system designed as a neural extension of SLPs is DeepStochLog (Winters et al. 2022). It is based on stochastic definite clause grammars, a type of SLPs, and introduces neural grammar rules. While DeepStochLog represents a step forward, it still faces challenges with scalability and expressiveness compared to DPrL. A more technical comparison between our method and DeepStochLog is reported in Section 7.
@@
Contributions. (i) We introduce DeepProofLog (DPrL), a novel deep SLP model that leverages neural parameterization at each proof step to provide continuous, goal-directed guidance during reasoning. (ii) We propose a theoretical mapping between derivation-based probabilistic logic reasoning and MDPs, enabling the use of powerful MDP solution techniques for efficient learning and inference in DPrL. (iii) Experiments on NeSy benchmarks and knowledge graph completion tasks empirically validate the effectiveness of our approach, achieving improved task performance, efficiency, and scalability over other NeSy systems.
@@ 2.1 Logic Programming
Following (Flach 1994; Lloyd 2012), we summarize the basic concepts of logic programming. A term is either a constant, a variable, or a compound term of the form $f(t_{1},...,t_{n})$, where $f$ is a functor of arity $n$ and each $t_{i}$ is a term. An expression $r(t_{1},...,t_{n})$, where $r$ is a predicate and each $t_{i}$ is a term, is called an atom. For example, $\textit{locatedIn}(X,\textit{europe})$ is an atom where europe denotes a constant and locatedIn a binary predicate. A definite clause [^2] We will omit “definite” as all clauses in this paper are definite. is of the form $H\leftarrow B_{1},...,B_{m}$, where $H$ is the head atom and $B_{1},\dots,B_{m}$ form the body atoms. A set of definite clauses $\mathcal{P}$ is called a definite logic program. A clause of the form $\leftarrow A_{1},\dots,A_{n}$ is called a goal and represents a query to the program. A substitution $\theta$ such as $\{X/italy\}$ is a mapping from variables to terms. A substitution $\theta$ is a unifier of two atoms (same for terms) $A_{1}$ and $A_{2}$ if $A_{1}\theta$ (application of $\theta$ to variables of $A_{1}$) = $A_{2}\theta$. For instance, $\theta=\{X/\textit{italy},Y/\textit{europe}\}$ unifies $A_{1}=\textit{locatedIn}(X,\textit{europe})$ and $A_{2}=\textit{locatedIn}(\textit{italy},Y)$, as $A_{1}\theta=A_{2}\theta=\textit{locatedIn}(\textit{italy},\textit{europe})$. A substitution $\theta$ is the Most General Unifier (MGU) if any other unifier can be obtained by further instantiating $\theta$—i.e., it makes the minimal necessary variable substitutions.
@@ 2.2 Stochastic SLD Resolution
Definition 1.
(Probability Distribution over SLD Derivations) Let $d=(G_{0},G_{1},\dots,G_{n})$ be a finite SLD derivation. The probability of $d$ is defined as:
@@
Let $\mathcal{R}(q)$ be the set of all successful derivations for a query $q$, i.e., derivations $d=(G_{0},G_{1},...,G_{n})$ with $G_{0}=q$ and $G_{n}=\textit{True}$, for some $n>0$. The success probability of $q$ is:
@@
DeepStochLog. DeepStochLog replaces fixed clause probabilities of SLPs with a neural network $\rho_{\lambda}$, which produces a probability distribution over possible variable substitutions. This induces a term-level neural parameterization of the transition model $p(G_{i+1}\!\mid\!G_{i})\!=\!\rho_{\lambda}(X\theta_{i},Y\theta_{i})$, where $\theta_{i}$ is the MGU used in the resolution step from $G_{i}$ to $G_{i+1}$, i.e., $\rho_{\lambda}$ is conditioned on the instantiated input and output terms $X\theta_{i},Y\theta_{i}$ realizing the unification step. An example is shown in Figure 1.
@@ 2.3 Markov Decision Processes
An MDP is a formal model of decision-making defined by the tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is a set of states, $\mathcal{A}(s)$ is a set of actions available in state $s$, $P(s,a,s^{\prime})$ is the transition function giving the probability of transitioning from state $s$ to $s^{\prime}$ with action $a$, $R(s,a,s^{\prime})$ is the reward function giving the expected reward for the transition, and $\gamma\in[0,1]$ is the discount factor for future rewards. A policy $\pi(a|s)$ defines the probability of selecting action $a$ in state $s$. Given a starting state $s_{0}$, solving an MDP requires finding the optimal policy $\pi^{*}$ that maximizes the future cumulative reward, defined as the value function:
@@ 3 DeepProofLog
To address the limitations of previous approaches, we propose a shift in perspective: instead of parameterizing the logic program itself (e.g., via clause weights), we parameterize the resolution process. To this end, we introduce DeepProofLog, which defines clause selection probabilities conditioned on the entire current goal and enables dynamic adaptation to varying substitution spaces. DeepProofLog is a framework for differentiable, goal-conditioned SLD resolution that supports flexible and context-aware reasoning.
@@
where $\mathcal{N}(G)$ denotes the set of next goals from resolving the leftmost atom of $G$; $f_{\lambda}(G,G^{\prime})$ scores the compatibility between the learned embeddings of the current goal $G$ and the candidate next goal $G^{\prime}$ (as defined in the next paragraph); and $\lambda$ comprises all model parameters.
@@
(Sub)symbolic Embeddings. Let $\mathcal{V}$ denote the set of all entities involved in reasoning, including symbolic tokens (e.g., predicates, constants, variables) and subsymbolic inputs (e.g., images). Each element $v\in\mathcal{V}$ is associated with a learnable dense embedding $\mathbf{e}_{v}\in\mathbb{R}^{d}$. The embeddings $e_{v}$ for all $v\in\mathcal{V}$ are part of $\lambda$ (cf. Equation 1) and are trained jointly with the reasoning model. This design enables a unified embedding space in which both symbolic and subsymbolic information can interact seamlessly during inference and learning.
@@
Compatibility Score. The compatibility score between goals $G$ and $G^{\prime}$ is computed as the scalar product of their embeddings: $f_{\lambda}(G,G^{\prime})=\mathbf{e}_{G}\cdot\mathbf{e}_{G^{\prime}}$.
@@
In line with similar approaches (cf. Section 2), DeepProofLog defines the query success probability as in Stochastic SLD resolution (cf. Definition 2). Moreover, our learning objective aligns with DeepStochLog,
@@ 4 Markov Decision Processes for DeepProofLog
DeepProofLog’s novel perspective on parameterizing stochastic SLD resolution opens the door to interpreting learning and reasoning in any (deep) SLP as learning an optimal policy within a logic program-based environment of an MDP. In the following we show how DeepProofLog instantiates an MDP by defining the components $(\mathcal{S},\mathcal{A},P,R,\gamma)$.
@@
Reward. To define the reward, we introduce an auxiliary set of labeled states, where each state is paired with the label of the original query it stems from, i.e., $\tilde{\mathcal{S}}=\{(G,y^{(i)})\mid G\in\mathcal{S},y^{(i)}\in\{0,1\}\text{ is the label of }q^{(i)}\in\mathcal{Q}\}$. Note that the labeled states $\tilde{\mathcal{S}}$ are only used for reward computation, do not affect the state space $\mathcal{S}$, and are not accessed by the policy. The reward function $R:\tilde{\mathcal{S}}\to\mathbb{R}$ is defined as
@@
As in standard MDPs, the value function of a policy is defined as its expected cumulative reward. While value functions generally follow the Bellman recursion, our SLD resolution MDP allows a simplified formulation for the value of a non-terminal goal $G$ with label $y$ (see Appendix A.2 for the derivation):
@@
We illustrate the introduced terms with the following example. Consider a logic program $\mathcal{P}=\{C_{1},C_{2},C_{3},C_{4},C_{5}\}$, where the clauses are defined as: $C_{1}=\textit{locIn}(X,Z)\leftarrow\textit{neighOf}(X,Y),\textit{locIn}(Y,Z)$, $C_{2}={}\leftarrow\textit{neighOf}(\textit{it},\textit{fr})$, $C_{3}={}\leftarrow\textit{locIn}(\textit{fr},\textit{eu})$, $C_{4}={}\leftarrow\textit{locIn}(\textit{tr},\textit{gr})$, and $C_{5}={}\leftarrow\textit{locIn}(\textit{gr},\textit{eu})$. Consider a positive query $q^{(0)}=G_{0}^{(0)}={}\leftarrow\textit{locIn}(\textit{it},\textit{eu})$ with label $y^{(0)}=1$. The corresponding action space is $\mathcal{A}(G_{0}^{(0)})=\{(C_{1},\{X/\textit{it},Z/\textit{eu}\})\}$. Applying the action gives the next goal $G_{1}^{(0)}={}\leftarrow\textit{neighOf}(\textit{it},Y),\textit{locIn}(Y,\textit{eu})$, with intermediate reward $R(G_{1}^{(0)})=0$. Assuming the episode ends after $n$ resolution steps with $G_{n}^{(0)}=\textit{True}$, the final reward is $R(G_{n}^{(0)},y^{(0)})=2\cdot 1-1=1$. Now consider a negative query $q^{(1)}=G_{0}^{(1)}={}\leftarrow\textit{locIn}(\textit{tr},\textit{eu})$ with label $y^{(1)}=0$. If the episode ends in $G_{m}^{(1)}=\textit{True}$ after $m$ resolution steps, the final reward is $R(G_{m}^{(1)},y^{(1)})=2\cdot 0-1=-1$.